# for Machine-Learned Interatomic Potentials

### v1

@@ 1 Introduction
Universal MLIPs—CHGNet1, MACE4, TensorNet27, ALIGNN35, SevenNet36, EquiformerV237—underpin high-throughput materials screening38, yet they are deployed without formal reliability guarantees. Matbench Discovery2 and earlier benchmarks10 evaluate 45+ models on 257K WBM (Wang–Botti–Marques3) structures, but aggregate accuracy metrics cannot answer the deployment-critical question: on which chemistries is this MLIP unreliable?
@@ 2 Related Work
Uncertainty quantification. Conformal prediction19 provides distribution-free marginal coverage guarantees applied to interatomic potentials20. Monte Carlo dropout29, deep ensembles30 and chemical-space uncertainty estimation31; 32 have been applied to molecular property prediction. Critically, these methods provide marginal coverage (“95% of all materials fall within the prediction interval”) but cannot answer the deployment question which specific compositions are unreliable and why. Conformal prediction intervals widen uniformly across compositional space; they do not identify the compositional drivers (e.g. heavy elements, large unit cells) that concentrate failures, nor do they enable targeted DFT allocation for high-risk chemistries. PCM concentrates adversarial sampling on failure boundaries, producing tighter envelopes where they matter most: in a split conformal baseline on the same 5,000 WBM materials (2,500 calibration, 2,500 test), the feature-space conformal envelope at contains 1,924 materials (precision 6.8%) versus PCM’s adversary-informed bootstrap envelope with 97 materials (precision 7.2%)—a reduction in envelope size at comparable precision. Active learning12; 50; 51 selects informative points for retraining; PCM selects adversarial points for auditing—the adversarial budget (200 queries) is orders of magnitude smaller than retraining data requirements.
Safety cases and formal verification. AMLAS16 provides structured safety-case methodology for ML components but relies on human argumentation. Seshia et al.17 identify the specification gap as a central challenge for verified AI. Neural network verification43; 44 focuses on input–output properties of classifiers; PCM verifies domain-specific reliability claims about regression models. Adversarial testing in ML33; 34 targets input perturbations; falsification of cyber-physical systems7; 8 provides compositional testing frameworks. PCM instantiates the Guaranteed Safe AI framework13 for materials: the MLIP is the world model, the envelope is the specification, and Lean 4 proofs are the verifier output—providing machine-checkable evidence that structured safety cases currently lack.
@@ 3.1 Architecture-specific failure profiles
We evaluate three architecturally distinct MLIPs—CHGNet v0.3.01, TensorNet (MatPES-PBE-v2025.1)27; 28, and MACE-MP medium4—on 5,000 WBM materials. Structures are synthetic compositional probes—atoms placed at irrational (golden-ratio) fractional coordinates to avoid symmetry artifacts, with DFT-derived cell volumes (Methods)—designed to test whether each MLIP has learned the correct energy landscape for a given chemistry. The blind-spot classification itself—which materials are DFT-stable but MLIP-unstable—derives from the WBM benchmark, where both DFT and MLIPs evaluate fully relaxed equilibrium structures. Force failure threshold: 50 eV Å-1 (Fig. 2).
@@ 4 Discussion
Formal verification and the specification gap. The Lean 4 proofs are explicitly conditional: every approximation (DFT at 0.1 eV atom-1, MLIP threshold at 2.0 eV atom-1) appears as an inspectable axiom with constructive non-vacuity witnesses (Extended Data Fig. Extended Data Fig. 9). This addresses the specification gap17 by making assumptions explicit rather than hiding them in implementation details. Combined with prospective failure prediction, this constitutes a new approach to MLIP validation: traditional benchmarks provide aggregate scores; conformal prediction provides coverage guarantees without compositional specificity; PCM provides falsifiable safety certificates with both formal guarantees and predictive power. Cross-MLIP transfer (AUC-ROC cross-model, feature importance ) confirms that vulnerability patterns are partially shared across architectures.
@@ 5.4 Lean 4 certification
Refined envelopes compile into five proof modules: core safety theorem, perturbation stability, error propagation (composing DFT + MLIP uncertainty via triangle inequality), monotone refinement soundness, and evaluation metric properties (CX count monotonicity, CI width convergence, PAC coverage guarantee). Lean 4.27.05, no mathlib dependency. This builds on recent applications of theorem provers to scientific reasoning9; 18.
@@ 5.8 Cross-MLIP evaluation
Three architecturally distinct MLIPs—CHGNet v0.3.01, TensorNet (MatPES-PBE-v2025.1-PES via matgl 2.0.6, evaluated in float32 precision), and MACE-MP medium4 (via mace-torch, float64)—are evaluated on 5,000 WBM materials (random sample, seed = 42). Structures are built from WBM composition, volume, and Wyckoff symmetry data using golden-ratio fractional coordinate spacing with correct DFT-derived cell volumes. Six pairwise correlations (energy and force) are computed with bootstrap 95% CIs (1,000 resamples). Force failure threshold: 50 eV Å-1.

### v2

@@ 1 Introduction
Universal MLIPs—CHGNet1, MACE4, TensorNet27, ALIGNN35, SevenNet36, EquiformerV237—underpin high-throughput materials screening38; 55, yet they are deployed without formal reliability guarantees59; 60. Matbench Discovery2, MLIP Arena57, CHIPS-FF58 and earlier benchmarks10 evaluate 45+ models on 257K WBM (Wang–Botti–Marques3) structures, but aggregate accuracy metrics cannot answer the deployment-critical question: on which chemistries is this MLIP unreliable?
@@ 2 Related Work
Uncertainty quantification. Conformal prediction19 provides distribution-free marginal coverage guarantees applied to interatomic potentials20. Monte Carlo dropout29, deep ensembles30 and chemical-space uncertainty estimation31; 32, committee disagreement63 and misspecification-aware regression61; 62 have been applied to molecular property prediction. Critically, these methods provide marginal coverage (“95% of all materials fall within the prediction interval”) but cannot answer the deployment question which specific compositions are unreliable and why. Conformal prediction intervals widen uniformly across compositional space; they do not identify the compositional drivers (e.g. heavy elements, large unit cells) that concentrate failures, nor do they enable targeted DFT allocation for high-risk chemistries. PCM concentrates adversarial sampling on failure boundaries, producing tighter envelopes where they matter most: in a split conformal baseline on the same 5,000 WBM materials (2,500 calibration, 2,500 test), the feature-space conformal envelope at contains 1,924 materials (precision 6.8%) versus PCM’s adversary-informed bootstrap envelope with 97 materials (precision 7.2%)—a reduction in envelope size at comparable precision. Active learning12; 50; 51 selects informative points for retraining; PCM selects adversarial points for auditing—the adversarial budget (200 queries) is orders of magnitude smaller than retraining data requirements.
Safety cases and formal verification. AMLAS16 provides structured safety-case methodology for ML components but relies on human argumentation. Seshia et al.17 identify the specification gap as a central challenge for verified AI. Neural network verification43; 44 focuses on input–output properties of classifiers; PCM verifies domain-specific reliability claims about regression models. Adversarial testing in ML33; 34 targets input perturbations; falsification of cyber-physical systems7; 8 provides compositional testing frameworks. Komendantskaya64 extends proof-carrying code to neuro-symbolic programs, and Kamran et al.65 apply it to LLM-generated code completions. PCM instantiates the Guaranteed Safe AI framework13 for materials: the MLIP is the world model, the envelope is the specification, and Lean 4 proofs are the verifier output—providing machine-checkable evidence that structured safety cases currently lack.
@@ 3.1 Architecture-specific failure profiles
We evaluate three architecturally distinct MLIPs—CHGNet v0.3.01, TensorNet (MatPES-PBE-v2025.1)27; 28, and MACE-MP medium4; 54—on 5,000 WBM materials. Structures are synthetic compositional probes—atoms placed at irrational (golden-ratio) fractional coordinates to avoid symmetry artifacts, with DFT-derived cell volumes (Methods)—designed to test whether each MLIP has learned the correct energy landscape for a given chemistry. The blind-spot classification itself—which materials are DFT-stable but MLIP-unstable—derives from the WBM benchmark, where both DFT and MLIPs evaluate fully relaxed equilibrium structures. Force failure threshold: 50 eV Å-1 (Fig. 2).
@@ 4 Discussion
Formal verification and the specification gap. The Lean 4 proofs are explicitly conditional: every approximation (DFT at 0.1 eV atom-1, MLIP threshold at 2.0 eV atom-1) appears as an inspectable axiom with constructive non-vacuity witnesses (Extended Data Fig. Extended Data Fig. 9). This addresses the specification gap17 by making assumptions explicit rather than hiding them in implementation details. Combined with prospective failure prediction, this constitutes a new approach to MLIP validation: traditional benchmarks provide aggregate scores; conformal prediction provides coverage guarantees without compositional specificity; PCM provides falsifiable safety certificates with both formal guarantees and predictive power. Cross-MLIP transfer (AUC-ROC cross-model, feature importance ) confirms that vulnerability patterns are partially shared across architectures. The recent extension of proof-carrying code to neuro-symbolic programs64 reinforces this direction: as ML models become embedded in scientific workflows, machine-checkable certificates become essential infrastructure, not optional overhead.
Emerging benchmarks and foundation models. Recent large-scale efforts—OMat2455 (118M DFT calculations), UMA56 (500M+ structures), and MatPES28—have dramatically expanded MLIP training data, while MLIP Arena57 and CHIPS-FF58 benchmark physics-aware properties beyond formation energy. These developments are complementary to PCM: larger training sets reduce but do not eliminate architecture-specific blind spots (as our cross-MLIP analysis with three models trained on overlapping Materials Project data demonstrates), and emerging benchmarks still evaluate aggregate accuracy rather than compositional specificity. PCM provides the missing layer—identifying which compositions remain unreliable after training on massive datasets, enabling targeted DFT validation where it matters most. Committee disagreement methods63 and misspecification-aware UQ61; 62 focus on per-prediction uncertainty but do not produce the explicit compositional failure maps or formal certificates that PCM provides. Robust training strategies66 and improved data generation67 address reliability from the training side; PCM addresses it from the deployment side—both are needed for trustworthy MLIP deployment. LLM-assisted materials design70 could further integrate with PCM by using adversarially discovered failure regions to guide generative exploration away from unreliable compositional space.
@@ 5.4 Lean 4 certification
Refined envelopes compile into five proof modules: core safety theorem, perturbation stability, error propagation (composing DFT + MLIP uncertainty via triangle inequality), monotone refinement soundness, and evaluation metric properties (CX count monotonicity, CI width convergence, PAC coverage guarantee). Lean 4.27.05, no mathlib dependency. This builds on recent applications of theorem provers to scientific reasoning9; 18; LLM-assisted proving68; 69 could further automate proof generation in future work.
@@ 5.8 Cross-MLIP evaluation
Three architecturally distinct MLIPs—CHGNet v0.3.01, TensorNet (MatPES-PBE-v2025.1-PES via matgl 2.0.6, evaluated in float32 precision), and MACE-MP medium4; 54 (via mace-torch, float64)—are evaluated on 5,000 WBM materials (random sample, seed = 42). Structures are built from WBM composition, volume, and Wyckoff symmetry data using golden-ratio fractional coordinate spacing with correct DFT-derived cell volumes. Six pairwise correlations (energy and force) are computed with bootstrap 95% CIs (1,000 resamples). Force failure threshold: 50 eV Å-1.